OpenStudy (anonymous):

Mr. Smith's art class took a bus trip to an art museum. The bus averaged 65 miles per hour on the highway and 25 miles per hour in a city. The art museum is 375 miles away from the school, and it took the class 7 hours to get there. Use Cramer's Rule to find how many hours the bus was on the highway and how many hours it was driving in the city.

OpenStudy (anonymous):

since no ones answering heres my answer..XD Let x be the number of miles traveled via highway Let y be the number of mile traveled via city routes Then: (1) x+y = 375 (2) x/65 + y/25 = 7 <==> (1/65)x + (1/25)y = 7 So our matrix looks like this: [ 1 1 ] [x] = [375] [ 1/65 1/25 ] [y] [ 7 ] (Below the |'s are determinants, as per Cramer's Rule) x = | 375 1 | | 7 1/25 | ------------------- | 1 1 | | 1/65 1/25 | = [(375*1/25)-(1*7)] / [1*1/25 - 1*1/65] = 8/ (8/325) = 325 miles y = | 1 375 | | 1/65 7 | ------------------- | 1 1 | | 1/65 1/25 | = [1*7 - (375*1/65)] / [1*1/25 - 1*1/65] = (16/13) / (8/325) = 50 miles So x = 325 miles. With average speed 65mi/h: 325mi/65mi/h = 5 hours y = 50 mi. At average speed 25mi/h: 50mi/25mi/h = 2 hours So, we travel 5 hours on the highway, and 2 hours on city streets.